OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -6 and 162, respectively.

OpenStudy (cwrw238):

nth term of a GS = a1r^(n-1) where r = common ration the r multiplies each term to get the next so 6th term / 2nd term = r^4

OpenStudy (cwrw238):

sorry its 2nd and 5th so we have 5th / 2nd = 162/ -6 = r^3

OpenStudy (cwrw238):

r^3 = -27 can you continue?

OpenStudy (anonymous):

are we jus cubing the numbers?

OpenStudy (cwrw238):

yea

OpenStudy (anonymous):

so r^3= 64

OpenStudy (cwrw238):

s r = cube root of -27

OpenStudy (anonymous):

so r= 3

OpenStudy (cwrw238):

now - its -27

OpenStudy (anonymous):

so -3

OpenStudy (cwrw238):

right now you can find the first term and plug that and r into the formula

OpenStudy (anonymous):

2 x (-3)^n-1

OpenStudy (cwrw238):

right

OpenStudy (cwrw238):

it might be better to leave the x out nth term = 2(-3)^(n-1)